# A Fluid-Structure Interaction Model of the Zebrafish Aortic Valve

### v1

@@ Abstract
The zebrafish is a valuable model organism for studying cardiac development and diseases due to its many shared aspects of genetics and anatomy with humans and ease of experimental manipulations. Computational fluid-structure interaction (FSI) simulations are an efficient and highly controllable means to study the function of cardiac valves in development and diseases. Due to their small scales, little is known about the mechanical properties of zebrafish cardiac valves, limiting existing computational studies of zebrafish valves and their interaction with blood. To circumvent these limitations, we took a largely first-principles approach called design-based elasticity that allows us to derive valve geometry, fiber orientation and material properties. In FSI simulations of an adult zebrafish aortic valve, these models produce realistic flow rates when driven by physiological pressures and demonstrate the spatiotemporal dynamics of valvular mechanical properties. These models can be used for future studies of zebrafish cardiac hemodynamics, development, and disease.
@@ 1 Introduction
Initially developed to model the human mitral valve (Kaiser et al.; 2019), we have previously applied design-based elasticity to the human aortic valve (Kaiser et al.; 2021). We studied the effect of phenotype on hemodynamics in Sievers Type 1 bicuspid valves (Kaiser et al.; 2022) and showed that model flows compared well to an in vitro experiment measured with flow with 4D flow Magnetic Resonance Imaging (4DMRI) (Kaiser et al.; 2023). We conducted initial simulation-guided design studies of bicuspidization repair of the aortic valve (Kaiser et al.; 2024) and validated predicted trends in pressure gradient in vitro (Choi, Sharir, Ono, Shibata, Kaiser, Zhu, Marsden, Woo, Ma and Kim; 2024; Choi, Sharir, Ono, Shibata, Kaiser, Palagani, Marsden and Ma; 2024). Our methods have thus proven effective for modeling a variety of flows involving heart valves and compare well to experimental data, and thus are appropriate for modeling the zebrafish valve.
Other FSI studies have examined flow in the developing heart (Battista et al.; 2018; Hsu et al.; 2019; Lee et al.; 2013, 2018; Vedula et al.; 2017), including one study with simplified two-dimensional endocardial cushions that develop into valves (Miller; 2011), and adult vasculature (Van Impe et al.; 2023).
To our knowledge this is the first study to perform three-dimensional FSI on the zebrafish valve. Our model produced realistic flow rates and robust closure in diastole over multiple cardiac cycles. This study serves as a proof of concept towards future studies of the mechanisms of zebrafish valve function, disease and development.
@@ 2.1 Imaging and Anatomy
Figure 1: Anatomy of zebrafish heart and light-sheet optical section of OFT (aortic) valves. A. Illustration of the adult zebrafish and its cardiac structure. B. Light-sheet imaging of a cleared 1 year-old zebrafish heart showing the AV and OFT valves.
@@ 2.2 Model construction
Models were constructed with a nearly first-principles method called design-based elasticity that we previously applied to model the human aortic valve (Kaiser et al.; 2021). The methods are briefly summarized here with changes for the zebrafish highlighted.
First, we assumed that tension in the leaflets supported a static, uniform pressure load $p$. The leaflet was represented as an unknown parametric surface $\mathbf{X}(u,v):\Omega\subset\mathbb{R}^{2}\to\mathbb{R}^{3}$, where curves of constant $v$ run circumferentially and curves of constant $u$ run radially. The leaflets exerted tension $S$ circumferentially, corresponding to the anatomical fiber direction, and tension $T$ radially, corresponding to the cross-fiber direction. We specified that tension on the boundary of an arbitrary patch of leaflet balanced pressure, applied the fundamental theorem of calculus to convert boundary terms to surface terms, then finally dropped the integrals as the region of interest was arbitrary. Let single bars, ||$\cdot$|| ⋅ |, denote the Euclidean norm and subscripts denote partial derivatives. We thus obtained the following system of partial differential equations that represent the predicted, loaded configuration of the leaflets,
@@
where $\alpha$ and $\beta$ control the maximum tension in each material direction, and $a$ and $b$ control spacing of the mathematical fibers in the model. Expression (2) acts, colloquially, as a “mathematical search function,” that allowed the solution to find a heterogenous field of tensions that support the specified load.
@@
To design a realistic model for a zebrafish, the pressure $p$ was set to 0.92 mmHg, the mean pressure difference across the valve when the aortic pressure was greater than the ventricular pressure (Hu et al.; 2001). The model was set to have two symmetrical leaflets, each subtending half the annular circumference. The valve diameter was set to 363 microns and commissure height to 285 microns, based on measurements from a light-sheet image of the valve (Fig. 1B). This imaging modality allowed viewing of the annulus, but was highly collapsed and compressed since the fish was deceased at the time of imaging. Thus, these are low estimates on the valve radius and height, but were used due to lack of comparable in vivo data. The parameters that control the maximum tension were set to $\alpha=4.1\cdot 10^{5}$ and $\beta=1.4\cdot 10^{4}$ dynes. The spacing parameter $a$ varied linearly from $a=26.2-100$ cm from the annulus to the free edge, and the parameter $b=40$ cm. These parameters were designed by trial and error to produce a model free edge length (leaflet width) of 571 microns in the reference configuration, or approximately 1.57 times the annular diameter, as in previous work on bicuspidization repair of the aortic valve (Kaiser et al.; 2024). The resultant leaflet height (or geometric height) was 169 microns in the reference configuration.
@@
To derive a constitutive law from the predicted, loaded configuration, we prescribed uniform stretch of $\lambda_{c}=1.15$ and $\lambda_{r}=1.54$ circumferentially and radially (Yap et al.; 2009). We derived the reference length $R$ for all links in the discretized model via solving the equation $\lambda=L/R$, where $\lambda$ is the relevant stretch and $L$ is the current length of a given edge or “link” in the discretized model. Each link was set to take an exponential response with exponential rate taken from experimental data (May-Newman et al.; 2009). The stiffness of each link was then scaled to take the tension required in the solution of the equilibrium equations (1) at the prescribed stretches. To obtain an open initial configuration, an additional equilibrium equation was solved with $p=0$ mmHg and the constitutive law just derived. Finally, two additional membranous layers were added to create a model thickness of 6.3 microns, which was scaled proportionally to valve radius from our previous models (Kaiser et al.; 2021). (Note that image resolution in the out of plane direction may cause the leaflets in Fig. 1B to appear artificially thicker than they are, and thus thickness was not assessed from this image.) The layers were attached to each other with stiff linear springs, and the stiffnesses of each membrane layer were set to one third of the initial stiffnesses across the three layers. The model was fully fiber-based, in that all circumferential, radial and cross-layer forces are computed along fibers, curves that exert tension along their axis. Links that were longer than twice the fluid mesh width were split and assigned the same relative spring constant. Stretch, stress and tangent modulus from FSI simulations were computed from the original membranous layer on the aortic side of the leaflet. This model construction resulted in a mean tangent modulus in the predicted loaded configuration of $2.83\cdot 10^{6}$ dynes/cm2 circumferentially and $1.16\cdot 10^{5}$ dynes/cm2 radially.
@@ 2.3 Fluid-Structure Interaction
Fluid-structure interaction simulations were performed with the Immersed Boundary Method (Peskin; 2002) in the open-source solver IBAMR (Immersed Boundary Adaptive Mesh Refinement) (Griffith et al.; 2010). The fluid mesh width was set to 5 microns. The time step was set to $5\cdot 10^{-7}$ s, which was extremely restrictive, due to very small spatial resolution and only modestly lower stiffness compared to human data. The valve was placed into a cylindrical tube of length 1.28 mm with diameter 363 microns with a flow extender at the inlet of length 150 microns and an additional radius 50 microns.
The cardiac cycle duration was 0.32 s, or approximately 185 beats per minute (Hu et al.; 2001) and simulations were run for two cycles. The first cycle was discarded due to initialization effects and results from the second cycle are shown. Times are reported from the start of the second cycle. The ventricular pressure was prescribed at the upstream inlet based on experimental measurements (Hu et al.; 2001), which were smoothed via convolution with a normalized cosine bump of radius .025 s and represented as a finite Fourier series with 600 frequencies.
The aortic outlet pressure was governed via a RCR (resistor capacitor resistor) lumped parameter network that models the distal vasculature, following a brief initialization phase. The minimum, maximum and mean aortic pressure were estimated to be 0.89, 2.08 and 1.39 mmHg respectively (Hu et al.; 2001). The flow rate target was set to 266 nl (Van Impe et al.; 2023). The ratio of proximal to distal resistors was set to 0.065 via an experimental estimate in humans, as we lack a value of this ratio for the zebrafish (Laskey et al.; 1990). The values were then estimated via the method previously described (Kaiser et al.; 2021). This process resulted in values $R_{p}=1.3\cdot 10^{4}$ s dynes cm-5, $R_{d}=2.1\cdot 10^{6}$ s dynes cm-5 and $C=1.0\cdot 10^{-7}$ cm5 dynes-1 for proximal resistance, distal resistance and capacitance, respectively.
@@ 3 Results
The valve shows unrestricted forward flow during systole and robust closure during diastole over multiple cardiac cycles (Fig. 2). The velocity field showed no evidence of regurgitation and low magnitude flow during closure. Twist was visible in the closed configuration, indicating presence of excess free edge length during diastole. The leaflets showed no billow or prolapse. The valve opens to nearly the radius of the cylindrical test chamber, indicating sufficient free edge length for maximal orifice area during systole. The forward flow is highly laminar, without any appearance of vortices or instabilities in the flow. When the flow rate nears zero, the local velocity is also low-magnitude while the valve remains open. The closing transient appears roughly symmetric, with a transient crease in both leaflets. Finally, the valve again fully closes with slight twist and no appearance of regurgitation. The flow dynamics are further depicted in the movie included with the supplemental information.
@@
Figure 2: Flows through the cardiac cycle, depicting a slice view of the vertical component of velocity $u_{z}$ with a side view of the valve (top row), the valve viewed en face (middle row) and the valve viewed at an angle (bottom row).
Pressure and flow rate waveforms showed approximately physiological values (Hu et al.; 2001; Van Impe et al.; 2023) through the cardiac cycle (Fig. 3). The aortic pressure resulted from the coupled dynamics of the full FSI system and the RCR boundary conditions. The pressure lowered gradually through end diastole, rose with systole, showed a prominent dicrotic notch upon valve closure, then again began to decrease gradually during diastole. The aortic pressure took minimum, maximum and mean values of 0.52, 2.14 and 1.10 mmHg, respectively. Across the valve (and including the surrounding test chamber), the peak pressure gradient was 0.62 mmHg and mean gradient was 0.34 mmHg. The flow rate was near zero during diastole, then showed rapid increase with forward pressure, followed by sustained forward flow. After the pressure difference changed sign, the flow rate shows a transient of backflow. The majority of this was likely blood that has not fully passed the leaflet free edge. The valve returned to its closed state, and the flow rate settled around zero without regurgitation.
The stroke volume was 266 nl and stroke volume minus retrograde volume (nearly all from the closing transient) was 209 nl. The maximum flow rate was $Q_{max}=$ 5.25 $\mu$l/s. The mean systolic flow was $Q_{mean}$ = 2.85 $\mu$l/s, which was computed between the time at which the pressure difference across the valve turned positive to the time at which the flow rate turned negative. The maximum Reynolds number was $Re_{max}=\rho(Q_{max}/A)r/\mu=2.30$ and the mean Reynolds number was $Re_{mean}=\rho(Q_{mean}/A)r/\mu=1.25$, where $A$ denotes the cross sectional area of the test chamber. The Reynolds number is thus order one and the flow lies in an intermediate regime, neither Stokes flow nor a fully inertial regime.
@@
The stretch, stress and tangent modulus were anisotropic and generally higher in diastole when the valve is closed (Fig. 4). In diastole, the circumferential stretch was near to the predicted loaded stretch, which is expected by construction, away from the free edge. The stretch was lower in the coapted region where the valve is presumably less loaded circumferentially. The stress and tangent modulus showed a concentration of stress and tangent modulus near the commissure. Stresses were more modest in the belly of the leaflet, and lower still at the free edge. The radial stretch, stress and tangent modulus were higher near the annulus and lower towards the free edge. The mean values of circumferential stretch, stress and tangent modulus were 1.12, $2.36\cdot 10^{4}$ dynes/cm2 and $1.36\cdot 10^{6}$ dynes/cm2, respectively. The mean values of radial stretch, stress and tangent modulus were 1.44, $4.09\cdot 10^{3}$ dynes/cm2 and $9.16\cdot 10^{4}$ dynes/cm2, respectively.

### v2

@@ Abstract
The zebrafish is a valuable model organism for studying cardiac development and diseases due to its many shared aspects of genetics and anatomy with humans and ease of experimental manipulations. Computational fluid-structure interaction (FSI) simulations are an efficient and highly controllable means to study the function of cardiac valves in development and diseases. Due to their small scales, little is known about the mechanical properties of zebrafish cardiac valves, limiting existing computational studies of zebrafish aortic valves and their interaction with blood. To circumvent these limitations, we took a largely first-principles approach called design-based elasticity that allows us to derive valve geometry, fiber orientation and material properties. In FSI simulations of an adult zebrafish aortic valve, these models produce realistic flow rates when driven by physiological pressures and demonstrate the spatiotemporal dynamics of valvular mechanical properties. These models can be used for future studies of zebrafish cardiac hemodynamics, development, and disease.
@@ 1 Introduction
Initially developed to model the human mitral valve (Kaiser et al.; 2019), we have previously applied design-based elasticity to the human aortic valve (Kaiser et al.; 2021). We studied the effect of phenotype on hemodynamics in Sievers Type 1 bicuspid valves (Kaiser et al.; 2022) and showed that model flows compared well to an in vitro experiment measured with flow with 4D flow Magnetic Resonance Imaging (4DMRI) (Kaiser et al.; 2023). We conducted initial simulation-guided design studies of bicuspidization repair of the aortic valve (Kaiser et al.; 2024) and validated predicted trends in pressure gradient in vitro (Choi et al.; 2024a, b). Our methods have thus proven effective for modeling a variety of flows involving heart valves and compare well to experimental data, and thus are appropriate for modeling the zebrafish aortic valve.
The zebrafish aortic valve has Reynolds number order 1 (Sec. 3), indicating an intermediate regime, neither stokes flow nor fully inertial. In contrast to the zebrafish, the human aortic valve operates in an inertial regime with Reynolds number $>10^{3}$ (Kaiser et al.; 2022). In this different flow environment, the human aortic valve has been extensively studied with FSI simulations. Many studies use the immersed boundary method (Peskin; 2002), as in this study, and its extensions (Kamensky et al.; 2015). Studies have examined jet analysis with Lagrangian coherent structures (Shadden et al.; 2010), hemodynamics associated with bicuspid aortic valve, (Marom et al.; 2013A; Emendi et al.; 2021; Kaiser et al.; 2022) and the interaction of aortic root and leaflet geometry in valve repair (Marom et al.; 2013B). Studies on prosthetic medical devices have modeled dynamics of valves in an in vitro pulse duplicator (Lee et al.; 2020), examined leaflet flutter (Lee et al.; 2021), revealed mechanisms of flow instabilities leading to turbulence (Bornemann and Obrist; 2024) and assessed hemodynamics for thrombosis potential (Bornemann et al.; 2024).
In the zebrafish, other FSI studies have examined flow in the developing heart (Battista et al.; 2018; Hsu et al.; 2019; Lee et al.; 2013, 2018; Vedula et al.; 2017), including one study with simplified two-dimensional endocardial cushions that develop into valves (Miller; 2011), and adult vasculature (Van Impe et al.; 2023).
To our knowledge this is the first study to perform three-dimensional FSI on the zebrafish aortic valve. Our model produced realistic flow rates and robust closure in diastole over two cardiac cycles. This study serves as a proof of concept towards future studies of the mechanisms of zebrafish aortic valve function, disease and development. Code for this project is open source and freely available (github.com/alexkaiser/heart_valves).
@@ 2.1 Imaging and Anatomy
Figure 1: Anatomy of zebrafish heart and light-sheet optical section of OFT (aortic) valves. A. Illustration of the adult zebrafish and its cardiac structure. B. Light-sheet imaging of a cleared 1 year of age zebrafish heart showing the AV and OFT leaflets.
@@ 2.2 Model construction
Models were constructed with a nearly first-principles method called design-based elasticity that we previously applied to model the human aortic valve (Kaiser et al.; 2021). In overview, to construct the model, we specified that tension in the leaflets must support a pressure and derived a resulting system of partial differential equations. Solving these equations yielded the predicted loaded configuration of the valve, with parameters and boundary conditions tuned to design the desired leaflet geometry. By temporarily prescribing a uniform pressure load in the model construction process, the necessary material properties to close under dynamic pressure loading in FSI simulations can be derived. Further, this method produces heterogeneous and anisotropic material properties and fiber orientations. This contrasts with standard techniques, in which material properties would be measured and prescribed from experimental data, which are highly challenging to obtain for a zebrafish.
The steps of this method are summarized as follows with changes for the zebrafish highlighted. First, we temporarily assumed that tension in the leaflets supported a static, uniform pressure load $p$. The leaflet was represented as an unknown parametric surface $\mathbf{X}(u,v):\Omega\subset\mathbb{R}^{2}\to\mathbb{R}^{3}$. Curves $\mathbf{X}(u,v=v_{0})$ on which $u$ varies for any given, fixed $v_{0}$ run circumferentially, corresponding to the anatomical fiber direction, and exerted tension $S$. Similarly, curves $\mathbf{X}(u=u_{0},v)$ on which $v$ varies for any fixed $u_{0}$ run radially, corresponding to the cross-fiber direction, and exerted tension $T$. We specified that tension on the boundary of an arbitrary patch of leaflet balanced pressure, applied the fundamental theorem of calculus to convert boundary terms to surface terms, then finally dropped the integrals as the region of interest was arbitrary. Let single bars, ||$\cdot$|| ⋅ |, denote the Euclidean norm and subscripts denote partial derivatives. We thus obtained the following system of partial differential equations that represent the predicted, loaded configuration of the leaflets,
@@
where $\alpha$ and $\beta$ control the maximum tension in each material direction, and $a$ and $b$ control spacing of the mathematical fibers in the model. Expression (2) acts, colloquially, as a “mathematical search function,” that allowed the solution to find a heterogeneous field of tensions that support the specified load. Non-uniform pressure or shear forces could be applied to the leaflets in this model construction step were their precise values available, but we expect the pressure to be close to uniform in FSI simulations and shear forces to be much smaller than pressure forces when the valve is closed. Note that in FSI simulations, the loading on the valve will include pressure, which is expected to be approximately uniform with minor spatial variation, and other fluid forces determined by their coupled interaction.
@@
To design a realistic model for a zebrafish, the pressure $p$ was set to 0.92 mmHg, the mean pressure difference across the valve when the aortic pressure was greater than the ventricular pressure (Hu et al.; 2001). The model was set to have two symmetrical leaflets, each subtending half the annular circumference. The valve diameter was set to 363 microns and commissure height to 285 microns, based on measurements from a light-sheet image of the valve (Fig. 1B). This imaging modality allowed viewing of the annulus, but was highly collapsed and compressed since the fish was deceased at the time of imaging. Thus, these are low estimates on the valve radius and height, but were used due to lack of comparable in vivo data. The parameters that control the maximum tension were set to $\alpha=4.1\cdot 10^{5}$ and $\beta=1.4\cdot 10^{4}$ dynes. The spacing parameters were taken to be dimensionless and $a$ varied linearly between $a=26.2-100$ from the annulus to the free edge, while $b=40$ throughout the leaflet. The leaflet mesh was constructed to have a mesh spacing of approximately 2.5 microns, half that of the fluid mesh. To achieve this, 384 points were placed along the annulus and 65 points in the radial direction. The resulting mesh had mean spacing 3.0 microns circumferentially and 1.6 microns radially at rest. The mesh spacing changed dynamically in the FSI simulations and was typically larger radially due to radial stretch through the cardiac cycle. These parameters were designed by trial and error to produce a model free edge length (leaflet width) of 571 microns in the reference configuration, or approximately 1.57 times the annular diameter, as in previous work on bicuspidization repair of the aortic valve (Kaiser et al.; 2024). The resultant leaflet height (or geometric height) was 169 microns in the reference configuration.
To derive a constitutive law from the predicted, loaded configuration, we prescribed uniform stretch of $\lambda_{c}=1.15$ and $\lambda_{r}=1.54$ circumferentially and radially (Yap et al.; 2009). We derived the reference length $R$ for all links in the discretized model via solving the equation $\lambda=L/R$, where $\lambda$ is the relevant stretch and $L$ is the current length of a given edge or “link” in the discretized model. The stretch-tension response of each link was modeled with an exponential function
$\tau=\kappa(e^{\eta(\lambda-1)}-1)$ (3)
with exponential rates $\eta$ = 57.46 circumferentially and $\eta$ = 22.40 radially based on experimental data (May-Newman et al.; 2009). The stiffness of each link $\kappa$ was then scaled to take the tension required in the solution of the equilibrium equations (1) at the prescribed stretches. This process resulted in anisotropic, heterogeneous material properties. The resultant mean tangent moduli circumferentially and radially were $2.83\cdot 10^{6}$ and $1.16\cdot 10^{5}$ dynes/cm2 respectively at the predicted loaded stretch ratios. The emergent stretches and tangent moduli in each direction were comparable but slightly lower in the simulation (See Sec. 3 and Fig. 5).
@@
To obtain an open initial configuration suitable for FSI simulations, an additional equilibrium equation was solved with $p=0$ mmHg and the constitutive law just derived. Finally, two additional membranous layers were added to create a leaflet thickness of 6.3 microns, which was scaled proportionally to valve radius from our previous models (Kaiser et al.; 2021). (Note that image resolution in the out of plane direction may cause the leaflets in Fig. 1B to appear artificially thicker than they are, and thus thickness was not assessed from this image.) This thickness was meant to model a nearly-zero-thickness, membranous structure. Previously, we found that a finite thickness of three layers spaced half the fluid mesh width apart, one fluid mesh width in total, mitigated the “grid-aligned artifact” that occurred when an immersed boundary supports a jump in pressure (Kaiser et al.; 2019). The layers were attached to each other with stiff linear springs with spring constant 78.90 dynes/cm, which was determined to be the stiffest value that did not result in further time step restrictions. The stiffnesses of each membrane layer were set to one third of the initial stiffnesses across the three layers. The model was fully fiber-based, in that all circumferential, radial and cross-layer forces are computed along fibers, curves that exert tension along their axis. Links that were longer than twice the fluid mesh width were split and assigned the same relative spring constant. Stretch, stress and tangent modulus from FSI simulations were computed from the original membranous layer on the aortic side of the leaflet.
@@ 2.3 Fluid-Structure Interaction
Fluid-structure interaction simulations were performed with the Immersed Boundary Method (Peskin; 2002) in the open-source solver IBAMR (Immersed Boundary Adaptive Mesh Refinement) (Griffith et al.; 2010). The fluid had a density of 1.0 g/ml and viscosity 0.04 poise. The spatial resolution was set to 5 microns and the time step was set to $5\cdot 10^{-7}$ s. The valve was placed in a cylindrical test chamber. Two cardiac cycles were simulated and results from the second cycle were analyzed. At the inlet, a pressure waveform from experimental measurements with heart rate 185 beats per minute was prescribed (Hu et al.; 2001). At the outlet, pressure boundary conditions were determined by an RCR (resistor capacitor resistor) lumped parameter network tuned for minimum, maximum and mean aortic pressure of 0.89, 2.08 and 1.39 mmHg respectively (Hu et al.; 2001) and stroke volume of 266 nl (Van Impe et al.; 2023). Additional details of the FSI setup and a mesh resolution study are shown in the Appendix.
@@ 3 Results
The valve shows unrestricted forward flow during systole and robust closure during diastole over two cardiac cycles (Fig. 2). The velocity field showed no evidence of regurgitation and low magnitude flow during closure. Twist was visible in the closed configuration, indicating presence of excess free edge length during diastole. The leaflets showed no billow or prolapse. The valve opens to nearly the radius of the cylindrical test chamber, indicating sufficient free edge length for maximal orifice area during systole. The forward flow is highly laminar, without any appearance of vortices or instabilities in the flow. When the flow rate nears zero, the local velocity is also low-magnitude while the valve remains open. The closing transient appears roughly symmetric, with a transient crease in both leaflets. Finally, the valve again fully closes with slight twist and no appearance of regurgitation. The flow dynamics are further depicted in the movie included with the supplemental information.
@@
Figure 2: Flows through the cardiac cycle, depicting a slice view of the vertical component of velocity $u_{z}$ with a side view of the valve (top row), the valve viewed en face with a slice view of the vertical component of velocity at the aortic annulus (middle row) and the valve viewed at an angle (bottom row). The penultimate frame shows the closing transient, followed by closure without regurgitation in the final frame.
Streamlines showed that the flow was laminar throughout the cardiac cycle (Fig. 3). Streamlines during forward flow showed no evidence of vortex shedding from the leaflets with very little tangential motion. When the valve was opening or closing, the streamlines near the leaflet went through the leaflet, since the leaflet was moving with the local fluid velocity. In the neighborhood of the leaflets, the streamlines showed some tangential components with the opening and closing leaflet kinematics, but remained largely vertical away from the valve. While the valve was closed, the streamlines appeared less regular due to remaining flows of near-zero magnitude. Streamlines during the closing transient also remained laminar, and at no time was a major vortex visible immediately on the aortic side of the leaflets.
Figure 3: Streamlines through the cardiac cycle. The streamlines are colored by the vertical component of velocity to illuminate the flow direction.
Pressure and flow rate waveforms showed approximately physiological values (Hu et al.; 2001; Van Impe et al.; 2023) through the cardiac cycle (Fig. 4). The aortic pressure resulted from the coupled dynamics of the full FSI system and the RCR boundary conditions. The pressure lowered gradually through end diastole, rose with systole, showed a prominent dicrotic notch upon valve closure, then again began to decrease gradually during diastole. The aortic pressure took minimum, maximum and mean values of 0.52, 2.14 and 1.10 mmHg, respectively. Across the valve (and including the surrounding test chamber), the peak pressure gradient was 0.62 mmHg and mean gradient was 0.34 mmHg. The flow rate was near zero during diastole, then showed rapid increase with forward pressure, followed by sustained forward flow. After the pressure difference changed sign, the valve initiated closure. During this time, the flow rate was transiently negative as the blood that was between the closing leaflets moved toward the ventricle. However, this does not indicate regurgitation. The valve returned to its closed state, and the flow rate settled around zero without regurgitation or leak.
The stroke volume was 266 nl and the maximum flow rate was $Q_{max}=$ 5.25 $\mu$l/s. The mean systolic flow was $Q_{mean}$ = 2.85 $\mu$l/s, which was computed between the time at which the pressure difference across the valve turned positive to the time at which the flow rate turned negative. The maximum Reynolds number was $Re_{max}=\rho(Q_{max}/A)r/\mu=2.30$ and the mean Reynolds number was $Re_{mean}=\rho(Q_{mean}/A)r/\mu=1.25$, where $A$ denotes the cross sectional area of the test chamber. The Reynolds number is thus order one and the flow lies in an intermediate regime, neither Stokes flow nor a fully inertial regime.
@@
The stretch, stress and tangent modulus were anisotropic and generally higher in diastole when the valve is closed (Fig. 5). In diastole, the circumferential stretch was near to the predicted loaded stretch, which is expected by construction, away from the free edge. The stretch was lower in the coapted region where the valve is presumably less loaded circumferentially. The stress and tangent modulus showed a concentration of stress and tangent modulus near the commissure. Stresses were more modest in the belly of the leaflet, and lower still at the free edge. The radial stretch, stress and tangent modulus were higher near the annulus and lower towards the free edge. The mean values of circumferential stretch, stress and tangent modulus were 1.12, $2.36\cdot 10^{4}$ dynes/cm2 and $1.36\cdot 10^{6}$ dynes/cm2, respectively. The mean values of radial stretch, stress and tangent modulus were 1.44, $4.09\cdot 10^{3}$ dynes/cm2 and $9.16\cdot 10^{4}$ dynes/cm2, respectively. The mean stretch in both directions was lower than the predicted stretch at which the models were designed, likely because the coaptation region contacts the opposing leaflet and is less loaded in this region. Accordingly, the mean tangent moduli were also modestly lower.